OpenStudy (anonymous):

integrate e^(x-1/x) / x^2 dx

OpenStudy (anonymous):

let x-1/x = u. Solving for du takes care of the denominator.

OpenStudy (raden):

first, simplify for (x-1)/x = x/x - 1/x = 1 - 1/x = 1 - x^(-1) u = 1 - x^(-1) du = x^(-2) dx or du = dx /x^2

OpenStudy (raden):

wait, which one ur problem :